### v3

@@ 3.3 Principal component analysis (PCA)
PCA is a statistical technique widely used for dimensionality reduction in data analysis. It transforms a high-dimensional dataset into a lower-dimensional set of linearly uncorrelated components that capture the maximum variance in the original data. In this study, we aim to understand patterns in human mobility data by applying PCA to the potential landscape of human flow. Our objective is to identify the positions of spatial grids within the potential landscape and examine how the principal components contribute to the temporal evolution of human flow patterns.
Consider a dataset represented by a matrix $\mathbf{X}$ with $n$ rows and $p$ columns. Each row $r$, denoted as $\mathbf{x}^{(r)}$, is a vector representing one observation of data with $p$ features. The data matrix $\mathbf{X}$ has column-wise zero empirical means, suggesting that the sample mean of each column has been shifted to zero. In this case, $\mathbf{X}$ represents the series of potentials across spatial grids, where $p$ is the number of spatial grids, and $n$ is the number of observations under various conditions. Specifically, $n=96$ captures data for 24 hours across different scenarios: weekdays in 2019, holidays in 2019, weekdays in 2021, and holidays in 2021. For example, a particular column may represent the potential at 08:00 on weekdays in 2019.
The primary goal of PCA is to identify features that explain most of the variance in the data. First, we calculate the covariance matrix $\mathbf{Q}$ of the dataset $\mathbf{X}$ as:
@@
Performing eigenvalue decomposition on $\mathbf{Q}$ yields a set of eigenvalues $\lambda_{1},\lambda_{2},\cdots,\lambda_{p}$ and their corresponding eigenvectors $\mathbf{w}_{1},\mathbf{w}_{2},\cdots,\mathbf{w}_{p}$. The eigenvectors $\mathbf{w}_{k}$ represent the directions in the feature space along which the data varies the most, while the eigenvalues $\lambda_{k}$ indicate the amount of variance captured by each corresponding eigenvector $\mathbf{w}_{k}$. The transformation is thereby performed as follows:
@@
which explicitly demonstrates PCA as a linear transformation that decomposes the mobility data into temporal components $\text{PC}_{k}^{(r)}$ and spatial patterns $\mathbf{w}_{k}$.
The $k$-th eigenvalue $\lambda_{k}$ is equal to the sum of the squares of the scores over the $n$ observations. It is expressed as:
@@
In this study, we analyze the eigenvectors and principal components scores to interpret the outcomes of PCA and their implications for human mobility patterns. As illustrated in Fig. 1 (right panel), PCA decomposes the potential landscape into distinct spatial and temporal dimensions. The axes PC1, PC2, and PC3 represent the spatial contributions of each principal direction, denoted by the eigenvector $\mathbf{w}_{k}$. The temporal evolution of mobility is captured through colored trajectories that depict the hourly variation in the principal component scores.
@@ 4.1 Temporal evolution of potential landscape
Using Japanese human mobility data (Sec. 2), we first calculated the potential landscape of human flow at different hours of the day. The results for weekdays in 2019 are presented in Fig. 2, displaying the spatial landscape at 3-hour intervals (hourly details are provided in the Supplementary Information). In these maps, the Yamanote railway loop is represented by a green line enclosing the Tokyo urban core, interlinking major centers such as Shinjuku, Shibuya, Ikebukuro, Ueno, and Tokyo Station. Areas beyond the loop primarily correspond to suburban regions and nearby cities.
During the morning rush hours, from 06:00 (Fig. 2c) to 09:00 (Fig. 2d), the potential values in the central area were markedly positive, whereas several grids in the peripheral areas exhibited negative potential values. This pattern reflected the significant movement of estimated commuters from residential zones to the central business district, where a high concentration of workplaces were located. The pronounced central attractiveness observed during these hours aligned with typical commuting patterns.
Figure 2: Time evolution of the potential landscape $-V(=s)$ for weekdays in 2019 in the Tokyo metropolitan area. Color codes represent the potential value in each spatial grid. Black contour lines delineate the borders of Tokyo’s special wards (Tokyo Metropolis), marking the primary urban center. Green contour lines denote the Yamanote railway loop, which encloses the city’s urban core. Subfigures correspond to different hours: (a) 00:00; (b) 03:00; (c) 06:00; (d) 09:00; (e) 12:00; (f) 15:00; (g) 18:00; (h) 21:00.
In the evening hours at approximately 18:00 (Fig. 2g), the peripheral areas exhibited distinctly positive potential values, while the urban core displayed negative values. This pattern marked the end of the workday and the outward movement of individuals from central workplaces as they returned to their residential areas. The trend persisted into the late evening (Fig. 2h), with negative potential values remaining in the central areas and positive values prevailing in the periphery. These trends also reflected post-work activities, such as social gatherings, before individuals returned home. These findings align with well-established commuting dynamics in urban environments, underscoring the utility of potential landscapes in capturing fundamental patterns of human mobility.
@@ 4.2 Spatiotemporal dynamics revealed through PCA on potential landscape
Further, we performed PCA on the potential landscapes for different scenarios, including weekdays and holidays in 2019 and 2021 (Sec. 3.3). Notably, the first three principal components accounted for 96.3% of the variance (Supplementary Information), underscoring their crucial role in explaining the underlying mobility patterns. By focusing on these dominant components, we transformed the temporal sequence of potential landscapes into a trajectory within the three-dimensional space of the principal components. As illustrated in the right panel of Fig. 1, this trajectory encoded temporal information, while the corresponding eigenvectors $\mathbf{w}_{k}$ conveyed the spatial characteristics of the flow. The results for weekdays in 2019 are shown in Fig. 3.
@@
Figure 3: Temporal evolution of the PC1, PC2, and PC3 scores, along with their corresponding eigenvectors on weekdays in 2019. (a) Time series of the principal component scores over 24-hour period. (b) and (c) Relationships between principal components in PC2 vs. PC1 and PC3 vs. PC1 spaces, respectively. The corresponding eigenvectors are illustrated in (d) $\mathbf{w}_{1}$, (e) $\mathbf{w}_{2}$, and (f) $\mathbf{w}_{3}$. Numbers in the maps (d)-(f) indicate key locations: 1-Yokohama, 2-Chiba, 3-Kawasaki, 4-Funabashi, 5-Ikebukuro, 6-Shinjuku, and 7-Tokyo station. Contour lines match those in Fig. 2.
@@
PC1 primarily captured the characteristic commuting patterns, with its score peaking during the morning rush hours (06:00-09:00) and dipping in the evening (17:00-22:00), as shown in Fig. 3a. The corresponding eigenvector $\mathbf{w}_{1}$ (Fig. 3d) was concentrated within the special wards of Tokyo, particularly inside the Yamanote loop line, which constituted the city’s central business district and extended to several key subcenters including Chiba, Kawasaki, and Yokohama. This spatial distribution, combined with the temporal trend, signified the typical flow of estimated commuters from suburban regions into urban centers in the morning, followed by return trips in the evening. Notably, PC1 consistently exhibited higher scores than PC2 and PC3, underscoring its primary role in capturing dominant mobility dynamics.
@@
PC2 reflected commuting behavior but exhibited more localized spatiotemporal characteristics. Its score peaked at approximately 08:00 and dropped between 17:00 and 18:00 (Fig. 3a). The eigenvector $\mathbf{w}_{2}$ (Fig. 3e) was positively concentrated around Tokyo Station and Shinagawa Station—both major high-speed rail hubs—and negatively concentrated around terminal stations such as Ikebukuro, Shinjuku, Kawasaki, Funabashi, and Yokohama. These terminals are positioned along multiple radial lines extending from Tokyo Station. Consequently, PC2 primarily reflected flows from these subcenter regions toward central high-speed rail hubs in the morning, followed by return trips in the early evening.
@@
Contrarily, PC3 exhibited a distinct temporal pattern, declining from 06:00 to 07:00 before sharply increasing between 08:00 and 10:00 (Fig. 3a). The eigenvector $\mathbf{w}_{3}$ (Fig. 3f) was predominantly positive in suburban areas and negative in central districts, with isolated positive hotspots near Tokyo and Yokohama Stations. These findings suggested that PC3 served as a complementary modulation to the commuting flows captured by PC1 and PC2, particularly reflecting movements toward or from key rail hubs and surrounding areas during morning hours. Furthermore, PC3 exhibited smoother yet more frequent variations throughout the day, capturing midday activity and smaller-scale movements, such as work-related travel, thereby contributing to the overall daily mobility patterns.
@@ 4.3 Comparison of human mobility patterns across scenarios
Figure 4: Comparison of weekday and holiday trajectories for 2019 in the principal component space. (a)-(c) Time evolution of PC1, PC2, and PC3 scores, respectively. (d) Trajectories projected in the PC1-PC2 plane. (e) Trajectories projected in the PC1-PC3 plane.
@@
The weekday and holiday score trajectories for 2019 are shown in Fig. 4, highlighting distinct temporal signatures. As discussed in Sec. 4.2, PC1 and PC3 captured weekday commuting flows and complementary movements. On holidays, PC1 followed a similar trajectory to weekdays, as shown in Fig. 4a, but with a lower overall amplitude and a peak occurring 1 hour later. The scores of PC3 exhibited a similar tendency, as shown in Fig. 4c. These findings indicated that commuting-like travel persisted on holidays but with reduced intensity and a later start compared to weekdays.
@@
PC2 exhibited markedly different behaviors on holidays than on weekdays. As shown in Fig. 4b, PC2 score was negative during holiday mornings (07:00-12:00) and positive in the evening (16:00-23:00), whereas on weekdays, PC2 score was positive in mornings and negative in the evening. This opposite trend in PC2 scores on holidays captured flows originating from central bullet train stations toward suburban terminals along radial lines, reversing the weekday flow pattern. Moreover, Figure 4b shows notable temporal differences in PC2 between weekdays and holidays. On weekdays, PC2 peaked sharply at approximately 08:00, while on holidays, it exhibited a broad dip from 07:00 to 12:00, reaching a minimum near 10:00. A similar shift was observed in the evening, with extended activity periods. These findings suggested that on holidays, PC2 primarily reflected leisure and entertainment trips rather than weekday commuting flows.
@@
Figure 5: Comparison of 2019 and 2021 trajectories for weekdays in the principal component space. (a)-(c) Time evolution of PC1, PC2, and PC3 scores, respectively. (d) Trajectories projected in the PC1-PC2 plane. (e) Trajectories projected in the PC1-PC3 plane.
@@
Further, we compared mobility trajectories before and during the COVID-19 pandemic. The weekday trajectories for 2019 and 2021 are shown in Fig. 5. Although both trajectories shared a similar overall shape—indicating that the fundamental structure of mobility patterns remained largely unchanged—the 2021 trajectory exhibited noticeably lower amplitudes during peak hours. For example, morning peaks in PC1 and PC2 were smaller (Fig. 5a-c), and the evening return-trip dip in PC2 was also reduced and ended earlier. Contrarily, the amplitude differences at other times of the day were less pronounced.
@@
During the data collection period (May and June 2021), the Japanese government declared a state of emergency in Tokyo to curb COVID-19 resurgence ahead of the 2021 Tokyo Summer Olympics. Although Japan’s mobility restrictions were relatively lenient compared to international measures, prior research documented an overall decline in metropolitan mobility [70, 71]. Our findings further revealed that these restrictions primarily reduced mobility during peak commuting hours, whereas daily movement patterns outside these periods remained relatively unchanged.

### v4

@@ 3.3 Principal component analysis (PCA)
PCA is a statistical technique commonly used for dimensionality reduction in data analysis. Previous research has used PCA to analyze and interpret human mobility patterns [SYW2011, CBT2013, SCL2021, LGL2021]. It transforms a high-dimensional dataset into a lower-dimensional set of orthogonal components that capture the maximum variance in the original data. In this study, we seek to characterize patterns in human mobility data by applying PCA to the potential landscape of human flow. Our objective is to identify the placement of spatial grids in the potential landscape and examine how the principal components contribute to the temporal evolution of human flow patterns.
Consider a dataset represented by a matrix $\mathbf{X}$ with $n$ rows and $p$ columns. Each row $r$, denoted as $\mathbf{x}^{(r)}$, is a vector representing a single observation with $p$ features. In this case, $\mathbf{X}$ represents a series of potentials across spatial grids, where $p$ is the number of spatial grids, and $n$ is the number of observations under various conditions. Specifically, $n=96$ covers 24 hours across different scenarios: weekdays in 2019, holidays in 2019, weekdays in 2021, and holidays in 2021. For example, a particular column might represent the potential at 08:00 on weekdays in 2019. $\mathbf{X}$ has zero column-wise means because the potentials given by Eq. 8 sum to zero, and the data were not standardized.
The primary goal of PCA is to identify features that explain the majority of the variance in the data. First, we compute the covariance matrix $\mathbf{Q}$ of the dataset $\mathbf{X}$ as:
@@
Performing eigenvalue decomposition on $\mathbf{Q}$ yields a set of eigenvalues $\lambda_{1},\lambda_{2},\cdots,\lambda_{p}$ and their corresponding eigenvectors $\mathbf{w}_{1},\mathbf{w}_{2},\cdots,\mathbf{w}_{p}$. The eigenvectors $\mathbf{w}_{k}$ represent the directions in the feature space along which the data vary the most. Meanwhile, the eigenvalues $\lambda_{k}$ indicate the amount of variance captured by each corresponding eigenvector $\mathbf{w}_{k}$. The transformation is then performed as follows:
@@
which clearly shows PCA as a linear transformation that decomposes the mobility data into temporal components $\text{PC}_{k}^{(r)}$ and spatial patterns $\mathbf{w}_{k}$.
The $k$-th eigenvalue $\lambda_{k}$ equals the sum of the squares of the scores over $n$ observations. It can be expressed as:
@@
In this study, we examine the eigenvectors and principal component scores to interpret the outcomes of PCA and their implications for human mobility patterns. As illustrated in Fig. 1 (right panel), PCA decomposes the potential landscape into separable spatial and temporal dimensions. The axes PC1, PC2, and PC3 represent the spatial contributions of each principal direction, denoted by the eigenvectors $\mathbf{w}_{k}$. The temporal evolution of mobility is captured through color-coded trajectories that depict the hourly variation in the principal component scores.
@@ 4.1 Temporal evolution of potential landscape
Using Japanese human mobility data (Section 2), we initially compute the potential landscape of human flow at different hours of the day. The results for weekdays in 2019 are presented in Fig. 4, displaying the spatial landscape at 3-h intervals (hourly details are provided in the Supplementary Information). In these maps, the Yamanote railway loop is represented by a green line enclosing the Tokyo urban core, linking major centers such as Shinjuku, Shibuya, Ikebukuro, Ueno, and Tokyo Station. Areas beyond the loop mainly comprise suburban regions and nearby cities.
During the morning rush hours, from 06:00 (Fig. 4c) to 09:00 (Fig. 4d), the potential values in the central area were strongly positive, whereas several grids in the peripheral areas exhibited negative potential values. This pattern indicated the significant movement of estimated commuters from residential zones to the central business district, where a high concentration of workplaces is located. The strong central pull observed during these hours aligned with typical commuting patterns.
Figure 4: Temporal evolution of the potential landscape V(=s)-V(=s) for weekdays in 2019 in the Tokyo metropolitan area. Grid colors represent the potential value in each spatial grid. Black contour lines delineate the borders of Tokyo’s special wards, marking the primary urban center. Green contour lines denote the Yamanote railway loop, which encloses the urban core of Tokyo. Subfigures how different hours: (a) 00:00; (b) 03:00; (c) 06:00; (d) 09:00; (e) 12:00; (f) 15:00; (g) 18:00; (h) 21:00.
In the evening hours at approximately 18:00 (Fig. 4g), the peripheral areas showed elevated levels, while the urban core displayed negative values. This pattern indicated the end of the workday and the outward movement of individuals from central workplaces as they returned to their residential areas. The trend persisted into the late evening (Fig. 4h), with negative potential values remaining in the central areas and positive values dominating in the periphery. These trends also captured post-work activities, such as social gatherings, before individuals returned home. These findings align with well-established commuting dynamics in urban environments, highlighting the utility of potential landscapes in capturing fundamental patterns of human mobility.
@@ 4.2 Spatiotemporal dynamics revealed through PCA on potential landscape
Additionally, we performed PCA on the potential landscapes for different scenarios, including weekdays and holidays in 2019 and 2021 (Section 3.3). Notably, the first three principal components accounted for 96.3% of the variance (Supplementary Information), highlighting their crucial role in explaining the underlying mobility patterns. By focusing on these dominant components, we mapped the temporal sequence of potential landscapes into a trajectory within the three-dimensional space of the principal components. As illustrated in the right panel of Fig. 1, this trajectory captured temporal information, while the corresponding eigenvectors $\mathbf{w}_{k}$ conveyed the spatial characteristics of the flow. These eigenvectors are the static potential landscapes, and their linear combinations with temporal coefficients approximate the temporal dynamics. The results for weekdays in 2019 appear in Fig. 5.
@@
Figure 5: Temporal evolution of the PC1, PC2, and PC3 scores, and their corresponding eigenvectors on weekdays in 2019. (a) Time series of the principal component scores over a day. (b) and (c) Relations between principal components in PC2 vs. PC1 and PC3 vs. PC1 spaces, respectively. The corresponding eigenvectors are shown in (d) $\mathbf{w}_{1}$, (e) $\mathbf{w}_{2}$, and (f) $\mathbf{w}_{3}$. Numbers on the maps (d)–(f) indicate key locations: 1-Yokohama, 2-Chiba, 3-Kawasaki, 4-Funabashi, 5-Ikebukuro, 6-Shinjuku, and 7-Tokyo Station. Contour lines mirror those in Fig. 4.
Figure 6: Correlations between the spatial characteristics of the principal components (eigenvectors $\mathbf{w}_{1}$, $\mathbf{w}_{2}$, $\mathbf{w}_{3}$) and spatial distribution of employees by industrial sectors.
@@
PC1 primarily captured the dominant commuting patterns. Temporally, its score peaks during the morning rush hours (06:00–09:00) and dips in the evening (17:00–22:00), as shown in Fig. 5a. Spatially, the potential landscapes from its corresponding eigenvector $\mathbf{w}_{1}$ (Fig. 5d) were concentrated within the special wards of Tokyo, particularly inside the Yamanote loop line, which constituted the central business district and extended to several key subcenters of the metropolis, including Chiba, Kawasaki, and Yokohama. This spatial distribution of potentials, combined with the temporal trend, indicated the typical flow of estimated commuters from suburban regions into urban centers in the morning, followed by return trips in the evening. Notably, PC1 consistently exhibited higher scores than PC2 and PC3, highlighting its primary role in capturing dominant mobility dynamics.
This interpretation finds support in the pearson-correlation analysis between the eigenvector $\mathbf{w}_{1}$ and the number of employees by industry (Fig. 6). The data on employees are drawn from the 2021 Economic Census for Business Activity conducted by the Statistics Bureau of Japan, and we used the major category defined by the Japan Standard Industrial Classification. Notably, the employees in the Agriculture and Forestry and Fisheries sectors are not individually shown in the data, and the Mining sector is omitted in this analysis because employment is minimal compared with other sectors. The eigenvector $\mathbf{w}_{1}$ of PC1 was positively correlated with the number of employees in every industry and showed a strong correlation (0.933 $\pm$ 0.007) with the total number of employees (all industries). This observation indicates that PC1 primarily captures the major commuting trips as the basic pattern of the daily mobility.
@@
PC2 captured commuting behavior but exhibited more localized spatiotemporal characteristics. Its score peaked at approximately 08:00 and dropped between 17:00 and 18:00 (Fig. 5a). The eigenvector $\mathbf{w}_{2}$ (Fig. 5e) was positive near Tokyo Station and Shinagawa Station—both major high-speed rail hubs—and negatively concentrated around terminal stations, such as Ikebukuro, Shinjuku, Kawasaki, Funabashi, and Yokohama. These terminals are located along multiple radial lines extending from Tokyo Station. As shown in Fig. 6, $\mathbf{w}_{2}$ of PC2 had a weak positive correlation ($\sim$0.25) with “G: Information and Communications,” whereas it shows a weak negative correlation ($\sim$ -0.25) with “N: Living-related and Amusement Services” and “M: Accommodations, Eating and Drinking Services.” This implies that PC2-driven flow is industry-related. Consequently, PC2 primarily captured industry-related flows from these subcenter regions toward central high-speed rail hubs in the morning, followed by return trips in the early evening.
@@
By contrast, PC3 exhibited a distinct temporal pattern, declining from 06:00 to 07:00 before sharply increasing between 08:00 and 10:00 (Fig. 5a). The eigenvector $\mathbf{w}_{3}$ (Fig. 5f) was predominantly positive in suburban areas and negative in central districts, with isolated positive hotspots near Tokyo and Yokohama Stations. $\mathbf{w}_{3}$ of PC3 had a weak positive correlation with “S: Government (except those classified elsewhere)” and “F: Electricity, Gas, Heat, and Water Supply,” whereas it shows a weak negative correlation with “O: Education and P: Medical, Health Care and Welfare.” These findings suggest that PC3 served as a complementary modulation to the commuting flows captured by PC1 and PC2. Early morning with a negative coefficient to $\mathbf{w}_{3}$, the employees in the education and medical sectors commute from suburban areas to central districts, and later. the employees in the government and public utility sectors move toward railway nodes. Furthermore, PC3 shows smoother yet more frequent variations throughout the day, capturing midday activity and smaller-scale movements, such as work-related travel, thereby contributing to the overall daily mobility patterns.
@@ 4.3 Comparison of human mobility patterns across scenarios
Figure 7: Comparison of weekday and holiday trajectories for 2019 in the principal component space. (a)–(c) Temporal evolution of PC1, PC2, and PC3 scores, respectively. (d) Trajectories projected onto the PC1-PC2 plane. (e) Trajectories projected onto the PC1-PC3 plane.
@@
The weekday and holiday score trajectories for 2019 appear in Fig. 7, highlighting distinct temporal signatures. As discussed in Section 4.2, PC1 and PC3 encode weekday commuting flows and complementary movements. On holidays, PC1 followed a similar trajectory to weekdays, as shown in Fig. 7a, but with a lower overall amplitude and a peak occurring 1 hour later. The scores of PC3 displayed a comparable pattern, as shown in Fig. 7c. These findings suggest that commuting-like travel persisted on holidays but with reduced intensity and a later start compared to weekdays.
@@
PC2 exhibited substantially different behaviors on holidays than on weekdays. As shown in Fig. 7b, the PC2 score was negative on holidays in the morning (07:00–12:00) and positive in the evening (16:00–23:00), whereas on weekdays, the PC2 score was positive in the morning and negative in the evening. This opposite trend in PC2 scores on holidays captured flows originating from central high-speed rail stations toward suburban terminals along radial lines, reversing the weekday flow pattern. As shown in Fig. 6, $\mathbf{w}_{2}$ of PC2 had a weak negative correlation ($\sim$ -0.25) with “N: Living-related and Amusement Services” and “M: Accommodations, Eating and Drinking Services.” Combined with its negative scores on holidays, this flow may represent leisure and entertainment trips. Moreover, Fig. 7b shows pronounced temporal differences in PC2 between weekdays and holidays. On weekdays, PC2 peaked sharply at approximately 08:00, while on holidays, it exhibited a broad dip from 07:00 to 12:00, reaching a minimum near 10:00. A similar shift appeared in the evening, with extended activity periods. These findings suggest that on holidays, PC2 primarily reflected leisure and entertainment trips rather than weekday commuting flows.
@@
Figure 8: Comparison of 2019 and 2021 trajectories on weekdays in the principal component space. (a)–(c) Temporal evolution of PC1, PC2, and PC3 scores, respectively. (d) Trajectories projected onto the PC1-PC2 plane. (e) Trajectories projected onto the PC1-PC3 plane.
@@
Additionally, we compared mobility trajectories before and during the COVID-19 pandemic. The weekday trajectories for 2019 and 2021 appear in Fig. 8. Although both trajectories had a comparable form—indicating that the fundamental structure of mobility patterns remained largely unchanged—the 2021 trajectory exhibited noticeably lower amplitudes during peak hours. For example, morning peaks in PC1 and PC2 were smaller (Fig. 8a–c), and the evening return-trip dip in PC2 was also reduced and ended sooner. By contrast, the amplitude differences at other times of the day were less pronounced.
@@
During the data collection period (May and June 2021), the Japanese government declared an emergency in Tokyo to curb the COVID-19 resurgence ahead of the 2021 Tokyo Summer Olympics. Although mobility restrictions in Japan were relatively lenient compared to international measures, prior research reported an overall decline in metropolitan mobility [KNR2022, HA2022]. Our findings Additionally show that these restrictions primarily reduced mobility during peak commuting hours, whereas daily movement patterns outside these periods remained relatively unchanged.